# Joint Transmission and Compression Design for 6G Networked Sensing with Limited-Capacity Backhaul

### v1

@@ IV-A Optimization of Transmit Covariance Matrices
Problem (P2.1) is a non-convex problem because $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$’s given in (II) are concave functions, instead of convex functions, of $\bm{R}_{n}$’s. We apply the successive convex approximation (SCA) technique to tackle this challenge. Specifically, because $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$ is a concave function, given any $\tilde{\bm{R}}_{n}\succeq\bm{0}$, $\forall n$, the following inequality holds
@@
$\mathbb{E}_{\bm{q}}\Bigg{\{}\text{tr}\Bigg{[}\bar{\bm{\Sigma}}_{n}^{-1}\Bigg{(}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}+\bar{\bm{Q}}_{n}\Bigg{)}\Bigg{]}\Bigg{\}}/\log 2$
$-M_{r}/\log 2+\mathbb{E}_{\bm{q}}\Big{[}\log_{2}|\bar{\bm{\Sigma}}_{n}|\Big{]}-\log_{2}|\bar{\bm{Q}}_{n}|,\quad\forall n,$ (33)
$\bar{\bm{\Sigma}}_{n}=\sum\limits_{u=1}^{N}\bm{G}_{n,u}\tilde{\bm{R}}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}_{n}+\bar{\bm{Q}}_{n}.$ (34)
Based on the above inequality, given any $\tilde{\bm{R}}_{n}\succeq\bm{0}$, $\forall n$, we aim to solve the following relaxed problem of Problem (P2.1)
@@
Note that in Problem (35), the non-convex constraint (29c) in Problem (P2.1) is replaced by the convex constraint (35b). Because Problem (35) is a convex problem, we can utilize the interior-point method to globally solve it.
@@ IV-B Optimization of Compression Noise Covariance Matrices
Problem (P2.2) is non-convex because $\bm{F}_{\boldsymbol{\xi}}(\{\bar{\bm{R}}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is not a concave function of $\bm{Q}_{n}$’s in (30b), and $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is not a convex function of $\bm{Q}_{n}$ in (29c). To tackle the above challenge, we first change the optimization variables. Specifically, define
@@
By replacing $\bm{Q}_{n}$ with $\bm{T}_{n}^{-1}-\sigma^{2}\bm{I}$, $\forall n$, in $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ that is characterized in Proposition 1, we can get $\bm{F}_{\boldsymbol{\xi}}(\{\bm{R}_{n},\bm{T}_{n}\}_{n=1}^{N})$. Moreover, the backhaul compression rate given in (II) becomes
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{T}_{n})=\log_{2}\Bigg{|}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\bm{T}_{n}^{-1}\Bigg{|}$
@@
Therefore, Problem (P2.2) is equivalent to
@@
s.t. $\begin{bmatrix}\bm{F}_{\boldsymbol{\xi}}(\{\bar{\bm{R}}_{n},\bm{T}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (39b)
@@
Now, (39b) is convex over $\bm{T}_{n}$’s. However, (39c) is still non-convex because $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{T}_{n})$ is a concave function, instead of convex function, of $\bm{T}_{n}$. Similar to the optimization of $\bm{R}_{n}$’s, we apply the SCA technique to tackle this challenge. Specifically, the relaxed problem of Problem (39) is
@@
$\mathbb{E}_{\bm{q}}\Big{\{}\text{tr}\left[\bar{\bm{\Omega}}_{n}^{-1}\Big{(}\bm{I}+\bar{\bm{J}}_{n}^{\frac{1}{2}}\bm{T}_{n}\bar{\bm{J}}_{n}^{\frac{1}{2}}\Big{)}\right]\Big{\}}/\log 2-M_{r}/\log 2$
@@
Problem (40) is a convex problem and can be solved by interior-point method. Then, the SCA-based algorithm to solve Problem (39) is summarized in Algorithm II. After the solution to Problem (39) is obtained, the compression noise covariance matrices can be obtained via (37).
@@
      Update $\tilde{\bm{T}}_{n}=\bm{T}^{(m-1)}_{n}$, $\forall n$, in Problem (40);
@@
V Numerical Results
In this section, we provide numerical examples to verify our results. The system consists of $N=2$ BSs, each equipped with $M_{t}=8$ transmit antennas and $M_{r}=8$ receive antennas, and $K=2$ targets. The locations of these two BSs are $\big{(}\frac{\sqrt{3}}{2},0\big{)}$ km and $\big{(}-\frac{\sqrt{3}}{2},0\big{)}$ km, respectively, and the two targets are randomly located at a small area around $\big{(}\frac{\sqrt{3}}{4},\frac{3}{4}\big{)}$ km and a small area around $\big{(}0,\frac{3}{4}\big{)}$ km, respectively. The power spectrum of the noise at the BSs is $-169$ dBm/Hz, and the channel bandwidth is $1$ MHz. We assume that the two BSs possess the same transmit power constraint, denoted by $\bar{P}=\bar{P}_{1}=\bar{P}_{2}$, and the same backhaul capacity constraint, denoted by $\bar{D}=\bar{D}_{1}=\bar{D}_{2}$.
To verify the effectiveness of Algorithm III to solve Problem (P2), we provide three benchmark schemes. Under Benchmark Scheme I, we perform uniform quantization to the received signals of each BS receive antenna as shown in [17] and merely optimize the transmit covariance matrices based on Algorithm I. Under Benchmark Scheme II, we fix the transmit covariance matrices of the BSs as $\bm{R}_{1}=\bm{R}_{2}=\frac{\bar{P}}{M_{t}}\bm{I}$ and merely optimize the compression noise covariance matrices based on Algorithm II. Under Benchmark Scheme III, we assume an ideal case where the backhaul capacity of each BS is infinite, i.e., constraint (29c) can be ignored in Problem (P2) and we merely design the transmit covariance matrices.

### v3

@@ IV-A Optimization of Transmission Covariance Matrices
Problem (P2.1) is a non-convex problem because $D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$’s given in (II) are concave functions, instead of convex functions, of $\bm{R}_{n}$’s. We apply the SCA technique to tackle this challenge, whose idea is to successively optimize the convex approximation of the original problem. Specifically, because $\hat{D}_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bar{\bm{Q}}_{n})$ is a concave function over the transmit covariance matrices $\{\bm{R}_{n}\}_{n=1}^{N}$, given any $\tilde{\bm{R}}_{n}\succeq 0,\forall n$, the following inequality always holds
@@
$\frac{1}{\log 2}\mathbb{E}_{\bm{q}}\Bigg{\{}\text{tr}\Bigg{[}\bar{\bm{\Sigma}}_{n}^{-1}\Bigg{(}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}+\bar{\bm{Q}}_{n}\Bigg{)}\Bigg{]}\Bigg{\}}$
$-\frac{M_{r}}{\log 2}+\mathbb{E}_{\bm{q}}\Big{[}\log_{2}|\bar{\bm{\Sigma}}_{n}|\Big{]}-\log_{2}|\bar{\bm{Q}}_{n}|,\quad\forall n,$ (36)
$\bar{\bm{\Sigma}}_{n}=\sum\limits_{u=1}^{N}\bm{G}_{n,u}\tilde{\bm{R}}_{u}\bm{G}_{n,u}^{H}+\sigma^{2}\bm{I}_{n}+\bar{\bm{Q}}_{n},\quad\forall n.$ (37)
Here, the first term at the right hand side of the equation (36) is a linear function on the transmit covariance matrix $\bm{R}_{n}$’s, and the other three terms are all constants. Based on the above inequality, given any $\tilde{\bm{R}}_{n}\succeq\bm{0}$, $\forall n$, we aim to solve the following relaxed problem of Problem (P2.1)
@@
Note that in Problem (38), the non-convex constraint (32c) in Problem (P2.1) is replaced by the convex constraint (38b). Because Problem (38) is a convex problem, we can utilize the interior-point method to globally solve it. Becauses of (35), the solution to Problem (38) is a feasible solution to Problem (P2.1).
@@ IV-B Optimization of Compression Noise Covariance Matrices
Problem (P2.2) is non-convex because $\bm{F}_{\bm{\xi}}(\{\bar{\bm{R}}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ is not a concave function of $\bm{Q}_{n}$’s in (56b), and $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{Q}_{n})$ is not a convex function of $\bm{Q}_{n}$ in (32c). To tackle the above challenge, we first change the optimization variables. Specifically, define
@@
By replacing $\bm{Q}_{n}$ with $\bm{T}_{n}^{-1}-\sigma^{2}\bm{I}$, $\forall n$, in $\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{Q}_{n}\}_{n=1}^{N})$ that is characterized in Proposition 1, we can get $\bm{F}_{\bm{\xi}}(\{\bm{R}_{n},\bm{T}_{n}\}_{n=1}^{N})$. Moreover, the fronthaul compression rate given in (II) becomes
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{T}_{n})=$ $\mathbb{E}_{\bm{q}}\left[\log_{2}\Bigg{|}\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{u}\bm{G}_{n,u}^{H}+\bm{T}_{n}^{-1}\Bigg{|}\right]$
@@
Therefore, if we change the design variables from $\bm{Q}_{n}$’s to $\bm{T}_{n}$’s, then Problem (P2.2) is equivalent to
@@
s.t. $\begin{bmatrix}\bm{F}_{\bm{\xi}}(\{\bar{\bm{R}}_{n},\bm{T}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (42b)
@@
Now, (42b) is convex over $\bm{T}_{n}$’s. However, (42c) is still non-convex because $D_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bm{T}_{n})$ is a concave function, instead of convex function, of $\bm{T}_{n}$. Similar to the optimization of $\bm{R}_{n}$’s, we can apply the SCA technique to tackle this challenge. Specifically, the relaxed problem of Problem (42) is
@@
$\frac{1}{\log 2}\mathbb{E}_{\bm{q}}\Big{\{}\text{tr}\left[\bar{\bm{\Omega}}_{n}^{-1}\Big{(}\bm{I}+\bar{\bm{J}}_{n}^{\frac{1}{2}}\bm{T}_{n}\bar{\bm{J}}_{n}^{\frac{1}{2}}\Big{)}\right]\Big{\}}-\frac{M_{r}}{\log 2}$
@@
Similar to (36), the function $\tilde{D}_{n}(\{\bar{\bm{R}}_{n}\}_{n=1}^{N},\bar{\bm{T}}_{n},\bm{T}_{n})$ is also a linear function over $\bm{T}_{n}$’s. Therefore, Problem (43) is a convex problem and can be solved by interior-point method. Then, the SCA-based algorithm to solve Problem (42) is summarized in Algorithm II. After the solution to Problem (42) is obtained, the compression noise covariance matrices can be obtained via (40).
@@
      Update $\bar{\bm{T}}_{n}=\bm{T}^{(m-1)}_{n}$, $\forall n$, in Problem (43);
@@
IV-C Computational Complexity Analysis
In this subsection, we provide the computational complexity analysis of the proposed algorithm for joint optimization of transmission covariance matrices and compression noise covariance matrices.
The computational complexity of proposed algorithm mainly comes from three parts. The first part is the derivation of the PFIM in (13) and the convex surrogate functions in (36) and (44), which consist of a lot of integral calculations. Let $\mathcal{O}(\Theta)$ denote the complexity of the integral of a function $f(\bm{q})$ over $\bm{q}$. Therefore, the complexity in this part can be given as $\mathcal{O}\left((4NK^{2}(M_{t}^{2}+M_{r}^{2})+2NI_{A}I_{S})\Theta\right)$ where $I_{A}$ and $I_{S}$ denotes the average number of iterations for the AO and the SCA-based optimization. Then we analyze the computational complexities of the other two parts of transmit covariance matrices optimization and the compression noise variance matrices optimization. For the former, the relaxed problem (38) in each iteration have $2(N+K)$ constraints, $M_{t}^{2}$ complex variables, and $2K$ real variables, the computational complexity can be given as $\mathcal{O}\left((2M_{t}^{2}+2K)^{3.5}+2(N+K)(2M_{t}^{2}+2K)^{2.5}\right)$. For the latter, in each iteration, the computational complexity of solving problem (43) is $\mathcal{O}\left((2M_{r}^{2}+2K)^{3.5}+2(N+K)(2M_{r}^{2}+2K)^{2.5}\right)$. At last, the whole computational complexity of Algorithm III can be given by $\mathcal{O}\big{(}4NK^{2}(M_{t}^{2}+M_{r}^{2})\Theta+I_{A}I_{S}(2N\Theta$ $+2(N+K)(2M_{t}^{2}+2K)^{2.5}+(2M_{t}^{2}+2K)^{3.5}+2(N+K)(2M_{r}^{2}+2K)^{2.5}+(2M_{r}^{2}+2K)^{3.5})\big{)}$.
V An Estimate-Then-Beamform-Then-Compress Strategy When $N_{r}$ is Large
In this section, we study how to reduce the complexity to design compression covariance matrices when the BS is equipped with a large number of receive antennas. A key observation from our numerical results (e.g., in Section VI) is that when $M_{r}$ is very large, many eigenvalues of the compression covariance matrices obtained via Algorithm III are extremely large. This indicates that when the dimension of the received signal is large, the BS will first transform the received signal into a low-dimension space and then perform quantization at these reduced dimensions. This is not surprising. When $M_{r}$ is large, if the quantization bits are allocated to quantize the signals received by all antennas, then the error of quantizing each signal is very large, leading to poor localization performance at the CU.
When $M_{r}$ is large, although at last only a low-dimension signal is quantized, at the beginning, we need to optimize the compression covariance matrices over a $M_{r}$-dimension space to find how to reduce the dimension of the received signal. As a result, as shown in Section IV-C, the complexity of Algorithm III is proportional to $M_{r}^{7}$. In this section, we will propose a low-complexity EBC strategy, under which the BS first estimates the AOAs of the signals from the targets (AOA estimation is quite accurate when $M_{r}$ is large), then beamforms the received signals into a low-dimension space, each pointing to one AOA direction, and last compresses the low-dimension signal and sends the quantization bits to the CU. Because compression design is over a lower dimension, the complexity of this scheme will be significantly reduced.
V-A Receive Beamforming and Compression Design
[…]
$\dot{\bm{y}}_{n}$ $=\bm{C}_{n}^{H}\bm{y}=\bm{C}_{n}^{H}\bm{A}_{n}\sum_{u=1}^{N}\bm{B}_{n,u}\bm{V}_{u}^{T}\bm{x}_{u}+\bm{C}_{n}^{H}\bm{z}_{n},~{}\forall n,$ (47)
where $\bm{C}_{n}\in\mathbb{C}^{M_{r}\times L_{r}}$ is the receive beamforming matrix at BS $n$. Then we also consider the Gaussian test channel for the compression process on $\dot{\bm{y}}_{n}$ as
$\ddot{\bm{y}}_{n}=\dot{\bm{y}}_{n}+\ddot{\bm{e}}_{n},\forall n,$ (48)
where $\ddot{\bm{e}}_{n}\sim\mathcal{CN}(\bm{0},\ddot{\bm{Q}}_{n})\in\bm{C}^{L_{r}\times 1}$ is the Gaussian quantization error vector with zero mean and covariance $\ddot{\bm{Q}}_{n}\in\mathbb{S}^{L_{r}}_{++}$ at BS $n$. Based on the signal model in (48), the fronthaul rate can be given by
$D_{n}(\{\bm{R}_{n}\}_{n=1}^{N},\bm{C}_{n},\ddot{\bm{Q}}_{n})=$
$\mathbb{E}_{\bm{q}}\left[\log_{2}\left|\bm{C}_{n}^{H}\left(\sum\limits_{u=1}^{N}\bm{G}_{n,u}\bm{R}_{n}\bm{G}_{n,u}^{H}\right)\bm{C}_{n}+\sigma^{2}\bm{I}+\ddot{\bm{Q}}_{n}\right|\right]$
$-\log_{2}\left|\ddot{\bm{Q}}_{n}\right|,~{}\forall n.$ (49)
Moreover, the PFIM $\ddot{\bm{F}}_{\bm{\xi}}$ can be modeled as
$\ddot{\bm{F}}_{\bm{\xi}}=\ddot{\bm{F}}_{0,\bm{\xi}}+\ddot{\bm{F}}_{p},$ (50)
where $\ddot{\bm{F}}_{p}$ is similarly to (III) with the refined target location distribution $\ddot{p}(\bm{q})$ based on the estimation, and the FIM is
$\ddot{\bm{F}}_{0,\bm{\zeta}}=\mathbb{E}_{\ddot{\bm{y}}}\left[\frac{\partial\log p(\ddot{\bm{y}}|\bm{\zeta})}{\partial\bm{\zeta}}\Bigg{(}\frac{\partial\log p(\ddot{\bm{y}}|\bm{\zeta})}{\partial\bm{\zeta}}\Bigg{)}^{T}\right],$ (51)
where $\ddot{\bm{y}}=[\ddot{\bm{y}}_{1},\dots,\ddot{\bm{y}}_{N}]^{T}$ is the aggregated received signals under the proposed EBC strategy. Based on the same approach used in Proposition 1, by simply changing $\bm{A}_{n}$ to $\bm{C}_{n}^{H}\bm{A}_{n}$, $\dot{\bm{A}}_{n}$ to $\bm{C}_{n}^{H}\dot{\bm{A}}_{n}$, and $\bm{O}_{n}$ to $\ddot{\bm{O}}_{n}=\sigma^{2}\bm{C}_{n}^{H}\bm{C}_{n}+\ddot{\bm{Q}}_{n}$, the FIM $\ddot{\bm{F}}_{0,\bm{\zeta}}$ can be obtained by the following proposition.
Proposition 2
The FIM $\ddot{\bm{F}}_{0,\bm{\zeta}}$ characterized in (51) can be expressed as
$\ddot{\bm{F}}_{0,\bm{\zeta}}=2\begin{bmatrix}\Re\{\ddot{\bm{F}}_{1}\}&\Re\{\ddot{\bm{F}}_{2}\}&-\Im\{\ddot{\bm{F}}_{2}\}\\ \Re\{\ddot{\bm{F}}_{2}\}^{T}&\Re\{\ddot{\bm{F}}_{3}\}&-\Im\{\ddot{\bm{F}}_{3}\}\\ -\Im\{\ddot{\bm{F}}_{2}\}^{T}&-\Im\{\ddot{\bm{F}}_{3}\}^{T}&\Re\{\ddot{\bm{F}}_{3}\}\end{bmatrix},$ (52)
where $\ddot{\bm{F}}_{1}$, $\ddot{\bm{F}}_{2}$, and $\ddot{\bm{F}}_{3}$ have the similar definition to (18) – (20), just with the terms $\left((\bm{A}_{i}^{l})^{H}\bm{O}^{-1}_{i}\bm{A}_{i}^{r}\right)$ and replaced by $\Big{(}(\bm{A}_{i}^{l})^{H}\bm{C}_{i}^{H}\ddot{\bm{O}}^{-1}_{i}\bm{C}_{i}\bm{A}_{i}^{r}\Big{)}$, for $i=1,\dots,N$, where $\bm{A}_{i}^{l},\bm{A}_{i}^{r}\in\Big{\{}\bm{A}_{i},\dot{\bm{A}}_{i}\Big{\}}$.
It is obvious that the receive beamforming matrices $\{\bm{C}_{n}\}_{n=1}^{N}$ also affect the FIM $\ddot{\bm{F}}_{0,\bm{\zeta}}$ via Proposition 2. In the special case with $\bm{C}_{n}=\bm{I},\forall n$, the FIM can be reduced to the one given by Proposition 1. Intuitively, utilizing $L_{r}<M_{r}$ beamformers at each BS can deteriorate the sensing performance even if the fronthaul capacity is unlimited. However, we can show that the estimation performance loss can be possibly avoided under appropriate receive beamforming design.
Theorem 1
In the case where the fronthaul capacity is unlimited such that there is no compression noise at each BS $n$, i.e., $\bm{Q}_{n}=\bm{0},\forall n$, there is no estimation performance loss if the receive beamforming matrix satisfies
$\mathcal{R}\{\bm{\Delta}_{n}(\bm{\theta}_{n})\}\subseteq\mathcal{R}\{\bm{C}_{n}\},\quad\forall n,$ (53)
where $\bm{\Delta}_{n}(\bm{\theta}_{n})=[\bm{A}_{n}(\bm{\theta}_{n}),\bm{\Pi}^{\perp}_{\bm{A}(\bm{\theta}_{n})}\dot{\bm{A}}(\bm{\theta}_{n})]\in\mathbb{C}^{M_{t}\times 2K}$ with $\bm{\Pi}^{\perp}_{\bm{A}(\bm{\theta}_{n})}=\bm{I}-\bm{A}(\bm{\theta}_{n})\left(\bm{A}(\bm{\theta}_{n})^{H}\bm{A}(\bm{\theta}_{n})\right)^{-1}\bm{A}(\bm{\theta}_{n})^{H}\in\mathbb{C}^{M_{t}\times M_{t}}$.
Proof:
Please refer to Appendix B
By Theorem 1, it illustrates that the receive beamforming is comprised by two parts of $\bm{A}_{n}(\bm{\theta}_{n})$ and $\bm{\Pi}^{\perp}_{\bm{A}(\bm{\theta}_{n})}\dot{\bm{A}}(\bm{\theta}_{n})$. The roles of these two parts can be intuitively understood as: beamformers in the column space of $\bm{\Pi}^{\perp}_{\bm{A}(\bm{\theta}_{n})}\dot{\bm{A}}(\bm{\theta}_{n})$ are essential to the target location determination in terms of AOA $\bm{\theta}_{n}$; while beamformers in the column space of $\bm{A}(\bm{\theta}_{n})$ are vital to preserve the information of the unknown attenuation coefficients $\bm{b}_{n}$, which also affects the localization performance.
Theorem 1 indicates that when $M_{r}$ is large, each BS can locally reduce the $M_{r}$-dimension signal to a $2K$-dimension signal without information loss in the ideal case with infinite fronthaul capacity. Note that the estimated AOAs $\hat{\bm{\theta}}_{n}$’s are very close to the true AOAs $\bm{\theta}_{n}$’s when $M_{r}$ is large. Therefore, we propose to design the beamforming vector of BS $n$ as
$\bm{C}_{n}=\bm{E}_{r}(\hat{\bm{\theta}}_{n})\in\mathbb{C}^{M_{r}\times 2K},~{}\forall n,$ (54)
where columns of $\bm{E}_{r}(\hat{\bm{\theta}}_{n})$ are the $L_{r}=2K$ eigenvectors corresponding to the $2K$ non-zeros eigenvalues of the matrix $\bm{\Delta}_{n}(\hat{\bm{\theta}}_{n})\bm{\Delta}_{n}^{H}(\hat{\bm{\theta}}_{n})$.
Then we can also formulate the joint transmission and compression covariance matrices optimization problem as
[…]
s.t. $~{}\text{tr}(\ddot{\bm{R}}_{n})\leq\bar{P}_{n},~{}\forall n,$ (55b)
$~{}D_{n}(\{\ddot{\bm{R}}_{n}\}_{n=1}^{N},\ddot{\bm{Q}}_{n})\leq\bar{D}_{n},~{}\forall n,$ (55c)
$~{}\ddot{\bm{Q}}_{n}\succeq\bm{0},~{}\forall n,$ (55d)
It is obvious that Problem (P3) is non-convex and we cannot directly solve it. However, we can easily find that Problem (P3) is very similar to Problem (P1), where only the dimension of variables are reduced. Thus, we can follow Section IV to solve Problem (P3). Specifically, we first employ Schur complement technique to transform Problem (P3) as
[…]
s.t. $~{}\begin{bmatrix}\ddot{\bm{F}}_{\bm{\xi}}(\{\ddot{\bm{R}}_{n},\ddot{\bm{Q}}_{n}\}_{n=1}^{N})&\bm{e}_{i}\\ \bm{e}_{i}^{T}&t_{i}\end{bmatrix}\succeq\bm{0},~{}\forall i,$ (56b)
[…]
Then, similarly, we also adopt the AO-based method to solve Problem (P4), which follows the procedure of Algorithm III to iteratively optimize $\ddot{\bm{R}}_{n}$’s and $\ddot{\bm{O}}_{n}$’s. It is also guaranteed that the objective function of Problem (P4) will be decreased after each iteration. Here, the two subproblems in each iteration of the AO-based method are solved by following Section IV-A and IV-B, respectively, whose details are omitted.
V-B Computational Complexity Analysis
Here, we also give the computational complexity for the joint transmission and compression optimization problem under the proposed hybrid estimation and compression strategy. Under the proposed EBC strategy, the subproblem for optimizing compression noise covariance matrices has $L_{r}^{2}$ complex variables and $2K$ real variables. Therefore, the complexity can be expressed as $\mathcal{O}\big{(}4NK^{2}(M_{t}^{2}+M_{r}^{2})\Theta+I_{A}I_{S}(2N\Theta+(2M_{t}^{2}+2K)^{3.5}+2(N+K)(2M_{t}^{2}+2K)^{2.5}+(2L_{r}+2K)^{3.5})+2(N+K)(2L_{r}+2K)^{2.5}\big{)}$. Since $L_{r}\ll M_{r}$ in the massive MIMO scenario, the computational cost can be greatly saved by adopting the proposed EBC strategy.
VI Numerical Results
In this section, we provide numerical examples to verify our results. The systems consists of $N=2$ BSs unless otherwise specified, whose locations are $\bm{\rho}_{1}=\left[\frac{\sqrt{3}}{2},0\right]^{T}$ km and $\bm{\rho}_{2}=\left[-\frac{\sqrt{3}}{2},0\right]^{T}$ km, respectively, and multiple targets whose location distributions follow
$p_{\bm{q}_{k}}(\bm{q}_{k})=\frac{1}{2\pi r_{k}^{2}}\exp\left(-\frac{||\bm{q}_{k}-\bar{\bm{q}}_{k}||_{2}^{2}}{2r_{k}^{2}}\right),~{}\forall k,$ (57)
where $\bar{\bm{q}}_{k}$ is the center of the possible location region of target $k$. The power spectrum of the noise at the BSs is $-169$ dBm/Hz, and the channel bandwidth is $1$ MHz. We assume that the two BSs possess the same transmit power constraint, denoted by $\bar{P}=\bar{P}_{1}=\bar{P}_{2}$, and the same fronthaul capacity constraint, denoted by $\bar{D}=\bar{D}_{1}=\bar{D}_{2}$.